you went to tae shop and found that the tae price is increased from rs.5 to rs.10. Therefore I, the demand for coffee increased from 20 cups to 50 cups. what will be the cross elasticity of demand

1
Expert's answer
2022-01-24T08:43:01-0500

"E=\frac{\frac{10}{5}}{\frac{50}{20}}=0.8" or 80%